OpenStudy (lovelyanna):

What is the value of the expression? 4!(7 – 4)!

OpenStudy (lovelyanna):

This is how I did it: 4*3*2*1 (7 - 4)! = 24 * 3! = 24 * (3*2*1) = 24 * 6 = 144 Final answer is 144, correct?

OpenStudy (anonymous):

You're right.

OpenStudy (lovelyanna):

Alright, thank you :)